A test was given to a group of students, their grades and sex are summarized below. A BC Total Male 8 6 10 24 Female 14 2 9 25 Total 22 8 19 49 If one student is chosen at random, find the probability that the student was NOT a (female that got a B). Preview Give your answer as a fraction or decimal out to at least 4 places.
Answer #1

Solution:- there are 2 females with B out of 49 students

probability that student is female with A = 2/49

probability not female with B = 1-(2/49) = 0.9592